# Vector Quantized-Elites: Unsupervised and Problem-Agnostic Quality-Diversity Optimization

### v2

@@ Abstract
Index Terms:
Quality-Diversity, Unsupervised Learning, MAP-Elites, Autonomous Robots
@@ I Introduction
Traditional evolutionary algorithms (EAs) have been a cornerstone of optimization, designed primarily to identify a single optimal solution for a given problem. Although effective for well-defined tasks, this narrow focus on singular objectives limits their ability to address complex challenges, such as those in robotics and autonomous systems. This limitation hinders their adaptability in dynamically changing environments, where a diverse range of high-performing solutions is often required. Consequently, traditional EAs struggle in scenarios that demand versatility rather than a single, best solution [1, 2].
Quality-Diversity (QD) algorithms address the limitations of traditional evolutionary algorithms by redefining the optimization objective [1, 2]. Rather than seeking a single optimal solution, QD algorithms, such as MAP-Elites [3, 4], aim to uncover a diverse set of high-performing solutions across a broad behavior landscape. This “illumination” approach generates multiple effective individuals, each associated with a unique behavior. By promoting both diversity and quality, QD algorithms enable more flexible and robust applications, particularly in autonomous systems, where a repertoire of high-quality behaviors is essential for adapting to dynamic challenges. For instance, QD algorithms have demonstrated the ability to generate diverse locomotion behaviors in quadruped robots [4, 5, 6].
@@
Early approaches to automatic BD extraction relied on large static datasets to learn lower-dimensional representations using techniques like Principal Component Analysis (PCA) or auto-encoders. For instance, Cully and Demiris [7] utilized a dataset of handwritten digits to train an auto-encoder, creating a low-dimensional latent space that captured the key features of different digits. This latent space then served as the behavior descriptor space, enabling QD algorithms to generate a diverse set of solutions spanning the learned feature space. This approach allowed a robotic arm to autonomously learn to draw digits without requiring manual definitions of their key characteristics.
The challenges posed by predefined BDs and static datasets encouraged the development of unsupervised learning methods for BD extraction. A notable example is the AURORA algorithm [8, 9], which employs neural network-based representation learning to autonomously identify behavior descriptors directly from the solution space. The main innovation compared to using a static dataset is that AURORA uses the data generated by the QD algorithm to train the auto-encoder in an online fashion. This innovation represented a significant leap, allowing QD algorithms to generalize across a broader range of tasks without relying on hand-crafted descriptors, thus enabling more adaptable and versatile QD frameworks. Nevertheless, AURORA still requires a lot of heuristics and extensive fine-tuning to be performant.
@@ II-A Quality-Diversity Algorithms
One of the notable strengths of QD algorithms is their ability to provide multiple high-performing solutions, offering users significant flexibility [1, 2]. This flexibility arises from the wide range of solutions available, allowing users to balance complexity, desired behavior, and practical implementation. Here, we present shortly the QD optimization framework for completeness, but we refer the reader to recent surveys for a more detailed overview [2, 1].
@@
The MAP-Elites algorithm [3] is one of the first and most popular QD algorithms. The popularity of MAP-Elites stems from its simplicity, both in terms of understanding and implementation. However, its main challenge lies in discretizing the behavior space, which is highly problem- and domain-dependent. This becomes particularly difficult in scenarios involving high-dimensional features or when the behavior space lacks well-defined constraints.
@@ II-B Unsupervised Representation Learning
VQ-VAE represents a powerful extension of traditional autoencoders, combining ideas from variational autoencoders (VAEs) and vector quantization. Introduced by Oord et al. [10], VQ-VAE addresses the limitations of continuous latent spaces by discretizing them, allowing the model to learn discrete latent representations that are often more interpretable and suitable for tasks such as generation, compression, and unsupervised learning.
@@
One of the key challenges of minimizing this combined loss is the non-differentiability of the quantization step (i.e., selecting the nearest codebook entry), which prevents gradients from flowing back through the encoder during standard training. To address this, VQ-VAE uses a straight-through estimator for gradient updates. Specifically, during the backward pass, the gradients from the decoder are passed through the quantization operation unchanged, which allows effective training despite the discrete nature of the latent space [10]. This paradigm shift to discrete latent spaces has led to breakthroughs in various fields, including natural language processing [11], computer vision [12, 13], speech recognition [14, 15], and reinforcement learning [16, 17].
@@ III Related Work
The MAP-Elites algorithm, introduced by Mouret and Clune [3], established a new paradigm in evolutionary optimization by illuminating diverse regions of the solution space rather than focusing solely on a single optimal solution. This class of algorithms later became known as Quality-Diversity algorithms [18]. MAP-Elites partitions the behavior space into user-defined niches and searches for the highest-performing solution within each niche. This approach has shown success in robotic control tasks, constrained optimization, and applications requiring broad behavioral diversity [4]. Other notable domains where MAP-Elites has been applied, showcasing its applicability and impact, are regarding the robustness of LLMs in adversarial prompts [19], game generation [20], evolving graph structures [21], and image synthesis [22]. However, MAP-Elites relies heavily on pre-defined behavior descriptors (BDs), which limits its flexibility and adaptability to dynamic or complex environments where optimal descriptors are not readily apparent.
@@
To address this limitation, various adaptations have been proposed. The CVT-MAP-Elites algorithm uses centroidal Voronoi tessellations (CVTs) to partition the behavior space based on well-spread centroids, allowing for more natural, high-dimensional behavior mappings without rigid grid structures [23]. Despite its effectiveness in certain applications, CVT-MAP-Elites still depends on predefined centroids, which may not capture the full range of possible behaviors in a dynamic environment.
@@
More recent work has explored integrating reinforcement learning (RL) techniques with MAP-Elites to improve search efficiency and expand behavior diversity. For example, methods like PGA-MAP-Elites and QD-PG leverage policy gradients alongside MAP-Elites’ mutation-based search, enabling gradient-informed exploration of high-dimensional behaviors [24, 25]. However, these approaches still rely on predefined descriptors or additional complexity from combining RL methods, which may not be ideal for fully unsupervised BD discovery.
@@
In parallel, advancements in skill discovery through hierarchical reinforcement learning (HRL) and skill chaining [26, 27] have shown the importance of decomposing tasks into reusable sub-skills. These methods provide a framework for discovering hierarchical behaviors but often require substantial supervision or domain knowledge. Unsupervised skill discovery methods, such as DIAYN (Diversity is All You Need) [28] and VALOR [29], emphasize maximizing diversity in the skill space while learning useful policies. However, these approaches primarily focus on policy-level diversity without leveraging structured archives like MAP-Elites for comprehensive exploration.
@@
The introduction of unsupervised learning methods into the QD framework, like AURORA [9], has enabled the automatic extracting latent behavior representations, allowing QD algorithms to operate independently of pre-defined descriptors. For instance, AURORA employs neural networks to identify meaningful patterns within the behavior space, yet it lacks a structured archive organization akin to MAP-Elites, which can hinder systematic exploration in complex domains [9] (as we will show in our experiments). The TAXONS algorithm [30], the predecessor of AURORA, also utilizes an autoencoder with RGB images as input to construct a low-dimensional behavior space, or outcome space as they characterize it. The authors, also use two metrics to decide the fitness of a solution, the fitness score and the autoencoder’s reconstruction error, and one of them is chosen for each iteration with a probability of 50%. More recently, STAX [31] operates through an alternating two-phase process: exploration, where it builds a diverse set of policies and learns the behavior representation; and exploitation, where it optimizes the performance of policies that have discovered rewards. Similarly, techniques leveraging deep reinforcement learning [32] and self-supervised approaches [33] have demonstrated promise in autonomous skill discovery, but their reliance on task-specific tuning can restrict generalization. Lastly, the Quality-Diversity Meta-Evolution (QD-Meta) framework [34], which evolves a population of QD algorithms to dynamically customize behavior spaces for user-defined meta-objectives. By using the Covariance Matrix Adaptation Evolutionary Strategy (CMA-ES), QD-Meta uses neural networks that transform high-dimensional behavior features into compact descriptors while leveraging reinforcement learning for dynamic parameter control. This approach enables the creation of tailored, diverse, and high-quality solutions across various domains without relying on manually crafted behavior spaces. Our approach, VQ-Elites, builds on these advancements by combining MAP-Elites with a VQ-VAE to autonomously generate structured, and flexible behavior spaces, thereby overcoming the limitations of previous unsupervised methods.
@@ IV Vector-Quantized Elites
The primary goal of this work is to advance the MAP-Elites to operate entirely unsupervised, eliminating the need for prior knowledge about the task, environment, or dynamic constraints, and in the process to propose a new paradigm for the development of unsupervised QD algorithms. Our proposed VQ-Elites algorithm achieves this by leveraging the agent’s experiences in conjunction with Vector Quantized (VQ) machine learning models. This combination allows VQ-Elites to dynamically and autonomously adapt to the specific characteristics of the problem at hand, ensuring flexibility and scalability across diverse tasks and environments. A detailed overview of the algorithm is provided in Algo. 1 and in Fig. 1.
@@
17:  $\mathcal{A}\leftarrow\mathcal{M}_{\text{encoder}}(\mathcal{D}_{\text{exploration}})$ $\triangleright$ Add random solutions to the archive
18:end procedure
19:procedure Evolution Phase
20:  $\boldsymbol{\theta}=$ selection($\mathcal{A}$) $\triangleright$ Select randomly solutions from the archive
21:  $\boldsymbol{\theta}^{\prime}=$ variation($\boldsymbol{\theta}$) $\triangleright$ Perform mutation and crossover
22:  $(fitness,d)\longleftarrow$ evaluate($\boldsymbol{\theta}^{\prime}$) $\triangleright$ Evaluate the solution, and get the fitness and the raw data
23:  $\mathcal{D}^{(i)}\longleftarrow d$ $\triangleright$ Add the raw data to the training dataset
24:  $\mathbf{bd}\longleftarrow\mathcal{M}_{encoder}(d)$ $\triangleright$ Extract the behavior descriptor
25:  $c\longleftarrow$ get_cell_index($\mathbf{bd}$, $\mathcal{C}$) $\triangleright$ Get the index of the cell in the grid the bd is assigned to
26:  if $\mathcal{A}(c)=null$ or $\mathcal{A}(c).fitness<fitness$ then $\triangleright$ Check if cooperation mode is on or if the cell is occupied and compete
27:   $\mathcal{A}(c)\longleftarrow fitness,\boldsymbol{\theta}^{\prime}$ $\triangleright$ Add solution to the archive
28:  end if
29:end procedure
30:procedure UpdateModel($\mathcal{D}$)
31:  train($\mathcal{M}$, $\mathcal{D}$) for $\mathcal{E}$ epochs $\triangleright$ Update the VQ-VAE model
32:  $\mathcal{C}\leftarrow\mathcal{M}_{\text{codebook}}$ $\triangleright$ Set the the updated codebook as the BD grid
33:  update_archive($\mathcal{A}$, $\mathcal{M}_{\text{encoder}}$) $\triangleright$ Update the BDs of the individuals in the archive, and reassign them to the closest cell
34:end procedure
A defining feature of the VQ-Elites algorithm is its unsupervised nature, which allows it to autonomously learn behavior descriptors for the discovered solutions. Unlike traditional QD algorithms, where users must manually define BDs [3, 6] or rely on prior knowledge to generate BDs using clustering methods [23], VQ-Elites removes these limitations. Manual or predefined BD selection can restrict the range of problems the algorithm can address or render it impractical in some cases. To overcome this, VQ-Eliteslearns BDs directly from the data.
VQ-VAE is a generative model that combines autoencoders with vector quantization to create discrete latent representations. VQ-VAEs learn a codebook of vectors, mapping continuous latent vectors onto discrete codes. For high-dimensional data such as images, videos, or multivariate time-series, this mapping typically occurs along the channel dimension. However, when applied to one-dimensional data, VQ-VAE performs gradient-based clustering of the latent vectors, where the discrete representations correspond to centroids — similar to the K-Means algorithm [35]. Hence, our algorithm can be used with any data modality, e.g., images, videos, or times-series. However, the latent extraction should be done by a fully connected layer on the flattened features extracted by the previous layers of the encoder.
@@
Alternative unsupervised methods for BD extraction, such as AURORA [9], face significant challenges when learning from online QD data. The dynamic nature of QD archives means that the behavior space evolves continuously, making it difficult for these methods to adapt effectively in real-time. Moreover, AURORA relies on additional hyperparameters, such as the minimum distance threshold for storing solutions in its unstructured archive. These hyperparameters can be difficult to tune and may require prior domain knowledge, reducing the method’s generality and scalability. While AURORA can be effective in certain scenarios, its reliance on unstructured archives and extra hyperparameters makes it less suitable for tasks that benefit from a structured, adaptive behavior space like the one provided by MAP-Elites.

### v3

@@ I Introduction
Traditional evolutionary algorithms (EAs) have been a cornerstone of optimization, designed primarily to identify a single optimal solution for a given problem. Although effective for well-defined tasks, this narrow focus on singular objectives limits their ability to address complex challenges, such as those in robotics and autonomous systems. This limitation hinders their adaptability in dynamically changing environments, where a diverse range of high-performing solutions is often required. Consequently, traditional EAs struggle in scenarios that demand versatility rather than a single, best solution [cully2017quality, chatzilygeroudis2021quality].
Quality-Diversity (QD) algorithms address the limitations of traditional evolutionary algorithms by redefining the optimization objective [cully2017quality, chatzilygeroudis2021quality]. Rather than seeking a single optimal solution, QD algorithms, such as MAP-Elites [mouret2015illuminatingsearchspacesmapping, cully2015robots], aim to uncover a diverse set of high-performing solutions across a broad behavior landscape. This “illumination” approach generates multiple effective individuals, each associated with a unique behavior. By promoting both diversity and quality, QD algorithms enable more flexible and robust applications, particularly in autonomous systems, where a repertoire of high-quality behaviors is essential for adapting to dynamic challenges. For instance, QD algorithms have demonstrated the ability to generate diverse locomotion behaviors in quadruped robots [cully2015robots, chatzilygeroudis2018reset, loco_srbd].
@@
Early approaches to automatic BD extraction relied on large static datasets to learn lower-dimensional representations using techniques like Principal Component Analysis (PCA) or auto-encoders. For instance, Cully and Demiris [cully2018hierarchical] utilized a dataset of handwritten digits to train an auto-encoder, creating a low-dimensional latent space that captured the key features of different digits. This latent space then served as the behavior descriptor space, enabling QD algorithms to generate a diverse set of solutions spanning the learned feature space. This approach allowed a robotic arm to autonomously learn to draw digits without requiring manual definitions of their key characteristics.
The challenges posed by predefined BDs and static datasets encouraged the development of unsupervised learning methods for BD extraction. A notable example is the AURORA algorithm [cully2019autonomous, aurora], which employs neural network-based representation learning to autonomously identify behavior descriptors directly from the solution space. The main innovation compared to using a static dataset is that AURORA uses the data generated by the QD algorithm to train the auto-encoder in an online fashion. This innovation represented a significant leap, allowing QD algorithms to generalize across a broader range of tasks without relying on hand-crafted descriptors, thus enabling more adaptable and versatile QD frameworks. Nevertheless, AURORA still requires a lot of heuristics and extensive fine-tuning to be performant.
@@ II-A Quality-Diversity Algorithms
One of the notable strengths of QD algorithms is their ability to provide multiple high-performing solutions, offering users significant flexibility [cully2017quality, chatzilygeroudis2021quality]. This flexibility arises from the wide range of solutions available, allowing users to balance complexity, desired behavior, and practical implementation. Here, we present shortly the QD optimization framework for completeness, but we refer the reader to recent surveys for a more detailed overview [chatzilygeroudis2021quality, cully2017quality].
@@
The MAP-Elites algorithm [mouret2015illuminatingsearchspacesmapping] is one of the first and most popular QD algorithms. The popularity of MAP-Elites stems from its simplicity, both in terms of understanding and implementation. However, its main challenge lies in discretizing the behavior space, which is highly problem- and domain-dependent. This becomes particularly difficult in scenarios involving high-dimensional features or when the behavior space lacks well-defined constraints.
@@ II-B Unsupervised Representation Learning
VQ-VAE represents a powerful extension of traditional autoencoders, combining ideas from variational autoencoders (VAEs) and vector quantization. Introduced by Oord et al. [vqvae_paper], VQ-VAE addresses the limitations of continuous latent spaces by discretizing them, allowing the model to learn discrete latent representations that are often more interpretable and suitable for tasks such as generation, compression, and unsupervised learning.
@@
One of the key challenges of minimizing this combined loss is the non-differentiability of the quantization step (i.e., selecting the nearest codebook entry), which prevents gradients from flowing back through the encoder during standard training. To address this, VQ-VAE uses a straight-through estimator for gradient updates. Specifically, during the backward pass, the gradients from the decoder are passed through the quantization operation unchanged, which allows effective training despite the discrete nature of the latent space [vqvae_paper]. This paradigm shift to discrete latent spaces has led to breakthroughs in various fields, including natural language processing [DBLP:journals/corr/abs-2312-11532], computer vision [vq-vae-2, walker2021predictingvideovqvae], speech recognition [DBLP:conf/icassp/WilliamsZCY21, DBLP:conf/icassp/ZhouBA21], and reinforcement learning [agrl, edl].
@@ III Related Work
The MAP-Elites algorithm, introduced by Mouret and Clune [mouret2015illuminatingsearchspacesmapping], established a new paradigm in evolutionary optimization by illuminating diverse regions of the solution space rather than focusing solely on a single optimal solution. This class of algorithms later became known as Quality-Diversity algorithms [pugh2016quality]. MAP-Elites partitions the feature space into user-defined niches and searches for the highest-performing solution within each niche. This approach has been utilized in multiple domains, showcasing its applicability and impact, regarding the robustness of LLMs in adversarial prompts [rainbow_teaming], game generation [gavel], evolving graph structures [nadizar2024searching], image synthesis [me_image_synthesis]. Additionally, MAP-Elites has been highly influential in robotics, specifically in robotic control tasks and constrained optimization, which require diverse behaviors (features) [cully2015robots]. However, MAP-Elites relies heavily on pre-defined behavior descriptors (BDs), which limits its flexibility and adaptability to dynamic or complex environments where optimal descriptors are not readily apparent.
@@
To address this limitation, various adaptations have been proposed. The CVT-MAP-Elites algorithm uses centroidal Voronoi tessellations (CVTs) to partition the behavior space based on well-spread centroids, allowing for more natural, high-dimensional behavior mappings without rigid grid structures [map-elites-voronoi]. Despite its effectiveness in certain applications, CVT-MAP-Elites still depends on predefined centroids, which may not capture the full range of possible behaviors in a dynamic environment.
@@
More recent work has explored integrating reinforcement learning (RL) techniques with MAP-Elites to improve search efficiency and expand behavior diversity. For example, methods like PGA-MAP-Elites and QD-PG leverage policy gradients alongside MAP-Elites’ mutation-based search, enabling gradient-informed exploration of high-dimensional behaviors [nilsson2021policy, pierrot2022qd]. However, these approaches still rely on predefined descriptors or additional complexity from combining RL methods, which may not be ideal for fully unsupervised BD discovery.
@@
In parallel, advancements in skill discovery through hierarchical reinforcement learning (HRL) and skill chaining [barto2003recent, konidaris2009skill] have shown the importance of decomposing tasks into reusable sub-skills. These methods provide a framework for discovering hierarchical behaviors but often require substantial supervision or domain knowledge. Unsupervised skill discovery methods, such as DIAYN (Diversity is All You Need) [eysenbach2018diversity] and VALOR [gregor2016variational], emphasize maximizing diversity in the skill space while learning useful policies. However, these approaches primarily focus on policy-level diversity without leveraging structured archives like MAP-Elites for comprehensive exploration.
@@
The introduction of unsupervised learning methods into the QD framework, like AURORA [aurora], has enabled the automatic extraction of latent behavior representations, allowing QD algorithms to operate independently of pre-defined descriptors. For instance, AURORA employs neural networks to identify meaningful patterns within the behavior space, yet it lacks a structured archive organization akin to MAP-Elites, which can hinder systematic exploration in complex domains [aurora] (as we will show in our experiments). The TAXONS algorithm [9196819], the predecessor of AURORA, also utilizes an autoencoder with RGB images as input to construct a low-dimensional behavior space, or outcome space as they characterize it. The authors, also use two metrics to decide the fitness of a solution, the fitness score and the autoencoder’s reconstruction error, and one of them is chosen for each iteration with a probability of 50%. More recently, STAX [paolo2024discovering] operates through an alternating two-phase process: exploration, where it builds a diverse set of policies and learns the behavior representation; and exploitation, where it optimizes the performance of policies that have discovered rewards. Similarly, techniques leveraging deep reinforcement learning [ecoffet2021go] and self-supervised approaches [gupta2018unsupervised] have demonstrated promise in autonomous skill discovery, but their reliance on task-specific tuning can restrict generalization. Lastly, the Quality-Diversity Meta-Evolution (QD-Meta) framework [9716036], which evolves a population of QD algorithms to dynamically customize behavior spaces for user-defined meta-objectives. By using the Covariance Matrix Adaptation Evolutionary Strategy (CMA-ES), QD-Meta uses neural networks that transform high-dimensional behavior features into compact descriptors while leveraging reinforcement learning for dynamic parameter control. This approach enables the creation of tailored, diverse, and high-quality solutions across various domains without relying on manually crafted behavior spaces. Our approach, VQ-Elites, builds on these advancements by combining MAP-Elites with a VQ-VAE to autonomously generate structured and flexible behavior spaces, thereby overcoming the limitations of previous unsupervised methods.
@@ IV Vector-Quantized Elites
The primary goal of this work is to advance the MAP-Elites to operate entirely unsupervised, eliminating the need for prior knowledge about the task, environment, or dynamic constraints, and in the process to propose a new paradigm for the development of unsupervised QD algorithms. Our proposed VQ-Elites algorithm achieves this by leveraging the agent’s experiences in conjunction with Vector Quantized machine learning models. This combination allows VQ-Elites to dynamically and autonomously adapt to the specific characteristics of the problem at hand, ensuring flexibility and scalability across diverse tasks and environments. A detailed overview of the algorithm is provided in Algo. 1 and in Fig. 1.
@@
17:end procedure
18:procedure Evolution Phase
19:  $\boldsymbol{\theta}=$ selection($\mathcal{A}$) $\triangleright$ Select randomly solutions from the archive
20:  $\boldsymbol{\theta}^{\prime}=$ variation($\boldsymbol{\theta}$) $\triangleright$ Perform mutation and crossover
21:  $(fitness,d)\longleftarrow$ evaluate($\boldsymbol{\theta}^{\prime}$) $\triangleright$ Evaluate the solution, and get the fitness and the raw data
22:  $\mathcal{D}^{(i)}\longleftarrow d$ $\triangleright$ Add the raw data to the training dataset
23:  $\mathbf{bd}\longleftarrow\mathcal{M}_{encoder}(d)$ $\triangleright$ Extract the behavior descriptor
24:  $c\longleftarrow$ get_cell_index($\mathbf{bd}$, $\mathcal{C}$) $\triangleright$ Get the index of the cell in the grid the bd is assigned to
25:  if $\mathcal{A}(c)=null$ or $\mathcal{A}(c).fitness<fitness$ then $\triangleright$ Check if cooperation mode is on or if the cell is occupied and compete
26:   $\mathcal{A}(c)\longleftarrow fitness,\boldsymbol{\theta}^{\prime}$ $\triangleright$ Add solution to the archive
27:  end if
28:end procedure
29:procedure UpdateModel($\mathcal{D}$)
30:  train($\mathcal{M}$, $\mathcal{D}$) for $\mathcal{E}$ epochs $\triangleright$ Update the VQ-VAE model
31:  $\mathcal{C}\leftarrow\mathcal{M}_{\text{codebook}}$ $\triangleright$ Set the the updated codebook as the BD grid
32:  update_archive($\mathcal{A}$, $\mathcal{M}_{\text{encoder}}$) $\triangleright$ Update the BDs of the individuals in the archive, and reassign them to the closest cell
33:end procedure
A defining feature of the VQ-Elites algorithm is its unsupervised nature, which allows it to autonomously learn behavior descriptors for the discovered solutions. Unlike traditional QD algorithms, where users must manually define BDs [mouret2015illuminatingsearchspacesmapping, loco_srbd] or rely on prior knowledge to generate BDs using clustering methods [map-elites-voronoi], VQ-Elites removes these limitations. Manual or predefined BD selection can restrict the range of problems the algorithm can address or render it impractical in some cases. To overcome this, VQ-Eliteslearns BDs directly from the data.
VQ-VAE is a generative model that combines autoencoders with vector quantization to create discrete latent representations. VQ-VAEs learn a codebook of vectors, mapping continuous latent vectors onto discrete codes. For high-dimensional data such as images, videos, or multivariate time-series, this mapping typically occurs along the channel dimension. However, when applied to one-dimensional data, VQ-VAE performs gradient-based clustering of the latent vectors, where the discrete representations correspond to centroids — similar to the K-Means algorithm [lancucki2020robust]. Hence, our algorithm can be used with any data modality, e.g., images, videos, or times-series. However, the latent extraction should be done by a fully connected layer on the flattened features extracted by the previous layers of the encoder.
@@
Alternative unsupervised methods for BD extraction, such as AURORA [aurora], face significant challenges when learning from online QD data. The dynamic nature of QD archives means that the behavior space evolves continuously, making it difficult for these methods to adapt effectively in real-time. Moreover, AURORA relies on additional hyperparameters, such as the minimum distance threshold for storing solutions in its unstructured archive. These hyperparameters can be difficult to tune and may require prior domain knowledge, reducing the method’s generality and scalability. While AURORA can be effective in certain scenarios, its reliance on unstructured archives and extra hyperparameters makes it less suitable for tasks that benefit from a structured, adaptive behavior space like the one provided by MAP-Elites.